OpenStudy (anonymous):

Write the equation of each parabola... a. with vertex (3,-1) and y-intercept 2 b. with vertex (0,5) passing through (2,9)

OpenStudy (anonymous):

Y=a(x-h)^2 + k This is the equation you have to plug in with your unknowns

OpenStudy (anonymous):

(h,k) is your vertex so those values are the following h= 3 k=-1

OpenStudy (anonymous):

so it'll be: 2 = a (0-3)^2 - 1 2 = a (-3)^2 -1 2= a (9) -1 2 = a + 8 4 = a?

OpenStudy (anonymous):

and the second one will be: y = a (x+0)^2 + 5 9 = a(2+0)^2 + 5 9 = a(2)^2 + 5 9 = a (4) + 5 9 = a (9) 1 = a?

OpenStudy (anonymous):

idk if it's right lol .____.

OpenStudy (anonymous):

The -1 doesn't subtract from the 9 because 9 has a coefficient of a so they are not like terms. You add 1 to both sides to get rid of the -1 . Then you divide by 9 to both sides

OpenStudy (anonymous):

That's the only mistake I saw but other than that you did very good.

OpenStudy (anonymous):

so for the first one, a = 1/3?

OpenStudy (anonymous):

Yes a=1/3 for the first one and you did the same mistake on the second one as well. You can't add the 5 and 4 together so you're going to have to subtract 5 to both sides. Then divide 4 to both sides but you still get the same answer as a=1 so that's weird lol

OpenStudy (anonymous):

oh LOL i see :) Thank you so much for your help! you're awesome xD